Find the functional values f(-5), f(1), and f(3) for the compound function.

f(x)={ |x-5|, if x is less than or equal to 1} { 1/(x+1), if x is greater than 1}
1
Expert's answer
2013-02-11T09:20:31-0500
f(x)={ |x-5|, if x is less than or equal to 1} { 1/(x+1), if x is greater than 1}
1) f(-5)
-5<1
Therefore:
f(-5) = | -5 - 5| = |-10| = 10
2) f(1)
1<=1
Therefore:
f(1) = |1-5| = |-4| = 4
3) f(3)
3>1
Therefore:
f(3) = 1/(3+1) = 1/4 = 0.25
Answer: f(-5) = 10, f(1) = 4, f(3) = 0.25.